OpenStudy (batgirl13):

PLEASE HELP ME ASAP!!!! WILL MEDAL AND FAN!!!! 15. The sun’s mass is about 2.7 x 107 times greater than the moon’s mass. The sun is about 400 times farther from Earth than the moon. How does the gravitational force exerted on Earth by the sun compare with the gravitational force exerted on Earth by the moon?

OpenStudy (batgirl13):

@matt101 @Mehek14 @pooja195

OpenStudy (astrophysics):

You may apply Newton's law of gravitation \[\huge F_g = \frac{ Gm_1m_2 }{ r^2 }\] where G is the gravitational constant of \[G = 6.67 \times 10^{-11}m^3kg^{-1}s^{-2}\] r = the distance between the centers of mass

OpenStudy (batgirl13):

So I would use the actual mass of the moon and sun? I know that sounds like a dumb question but this question is driving me insane

OpenStudy (astrophysics):

Well we want to solve for a ratio, so you can find it separately by using all the quantities such as moons mass etc for the force exerted on earth by moon and the sun. Then setting it up as a ratio, but we have no need for that. We can solve for \[\frac{ F_s }{ F_m }\] where the subscripts represent gravitational force of sun and m is the gravitational force of the moon. There will be lots of cancellations when you plug in the formula, see if you can do that. And if you get stuck, don't worry we can go over it together.